# A Reinforcement Learning Framework for Application-Specific TCP Congestion-Control

### v1

@@ IV Experiments
In this section, we extensively evaluate ASC from five perspectives: 1) achieving the optimization objective, 2) impact and overhead of the updating interval, 3) robustness in dynamically changing environments, 4) effectiveness of FRM in adapting to new objectives, and 5) fairness in utilizing network resources. We conducted our evaluations on three networks with well-known topologies [39], namely dumbbell, leaf-spine, and fat-tree, shown in Fig. 3. The bottleneck bandwidths in each topology were kept at 2Mbps. All other links had the capacity of 10Mbps.
In addition to measuring the performance of ASC, we further compared it with prior approaches. For this, we selected two recent DRL-based algorithms, TCPAI [40] and PCCRL [24] as well as two heuristic-based algorithms, NewReno [16] and Cubic [2]. We emphasize here that while we compare ASC with TCPAI and PCCRL, both TCPAI and PCCRL suffer from all the limitations mentioned in Sec. I-B. In particular, note that they can only achieve a single predefined optimization goal, i.e., I-Apps cannot set their own customized optimization objectives. Next, we present observations from our experiments conducted in ns-3 [41] on the five perspectives listed above.
@@ IV-A Achieving the Objective
Figs. 4(b) and 4(c) show latency and jitter, respectively, experienced by the five approaches in each of our three topologies when the objective we set for ASC was to minimize latency and to minimize jitter, respectively. Again, ASC outperforms all other approaches, except CUBIC when optimizing latency under Leaf-Spine and Fat-Tree topologies. However, note from Fig. 4(a) that CUBIC achieved the lowest throughput, i.e., it keeps such little data in-flight that its packets experience hardly any congestion, and thus see the lowest latency. While we saw in Fig. 4(a) that the throughput’s achieved by some of the other approaches were close to ASC’s, we see a significantly better performance achieved by ASC in latency and jitter compared to the other approaches. In fact, for New Reno and Cubic, we had to employ a secondary y-axis due to their very high jitter compared to ASC. This significant performance difference between ASC and prior approaches is because prior approaches could only handle one network metric. Most prior approaches focused on optimizing just the throughput despite extensive need by various I-Apps for optimizing other metrics, such as latency, jitter, packet loss, and so on. This experimentally demonstrates the flexibility that our approach offers to network administrators as well as I-App developers in setting diverse optimization objectives for their I-Apps.
@@ V Conclusion
In this paper, we presented ASC, a DRL based framework for TCP congestion control that separates the deployment of learning and inference modules, making it computationally very simple for use on commodity devices. The architecture of ASC inherently makes it highly scalable. They key technical novelty of ASC lies in its ability to handle arbitrary optimization objectives from I-Apps as well as its ability to handle changes in objectives of I-Apps at runtime. The key technical depth of ASC lies in its design of the reward function that not only ensures that I-Apps achieve their desired objectives but also ensures that the traffic of the I-Apps traversing the network demonstrates fair use of the available resources. We have demonstrated these aspects extensively through experiments. Our results further showed that ASC can not only achieve arbitrary objectives specified by I-Apps, it actually outperformed all prior approaches even in the specific objectives they were designed to achieve. We envision that the framework introduced in this paper will serve as the foundational stepping stone for the research community to build further on it and harness the potential that the learning-based approaches hold for TCP congestion control.

### v2

@@ IV Experiments
In this section, we extensively evaluate ASC from five perspectives: 1) achieving the optimization objective, 2) impact and overhead of the updating interval, 3) robustness in dynamically changing environments, 4) effectiveness of FRM in adapting to new objectives, and 5) fairness in utilizing network resources. We conducted our evaluations on three networks with well-known topologies [42], namely dumbbell, leaf-spine, and fat-tree, shown in Fig. 3. The bandwidth of bottleneck links (thick black) in each topology was kept at 2Mbps. All other links had the capacity of 10Mbps. Each simulation experiment ran for 60,000 steps, each step spanning 50ms. For every configuration, we conducted 10 independent runs and report aggregate statistics. Application traffic followed periodic ON/OFF burst models. Unless otherwise specified, the learning rate is set to 1e-4 and SAC is trained with a buffer size of 1e5.
In addition to measuring the performance of ASC, we further compared it with prior approaches. For this, we selected two recent DRL-based algorithms, TCPAI [43] and PCCRL [19] as well as two heuristic-based algorithms, NewReno [44] and Cubic [2]. We emphasize here that while we compare ASC with TCPAI and PCCRL, both TCPAI and PCCRL suffer from all the limitations mentioned in Sec. I-A. In particular, note that they can only achieve a single predefined optimization goal, i.e., I-Apps cannot set their own customized optimization objectives. Next, we present observations from our experiments conducted in ns-3 [45] on the five perspectives listed above.
@@ IV-A Achieving the Objective
Figs. 4(b) and 4(c) show latency and jitter, respectively, experienced by the five approaches in each of our three topologies when the objective we set for ASC was to minimize latency and to minimize jitter, respectively. Again, ASC outperforms all other approaches, except CUBIC when optimizing latency under Leaf-Spine and Fat-Tree topologies. However, note from Fig. 4(a) that CUBIC achieved the lowest throughput, i.e., it keeps such little data in-flight that its packets experience hardly any congestion, and thus see the lowest latency. While we saw in Fig. 4(a) that the throughput’s achieved by some of the other approaches were close to ASC’s, we see a significantly better performance achieved by ASC in latency and jitter compared to the other approaches. In fact, for New Reno and Cubic, we had to employ a secondary y-axis due to their very high jitter compared to ASC.
@@ V Conclusion
In this paper, we presented ASC, a DRL based framework for TCP congestion control that separates the deployment of learning and inference modules, making it computationally very simple for use on commodity devices. The architecture of ASC inherently makes it highly scalable. They key technical novelty of ASC lies in its ability to handle arbitrary optimization objectives from I-Apps as well as its ability to handle changes in objectives of I-Apps at runtime. The key technical depth of ASC lies in its design of the reward function that not only ensures that I-Apps achieve their desired objectives but also ensures that the traffic of the I-Apps traversing the network demonstrates fair use of the available resources. We have demonstrated these aspects extensively through experiments. Our results further showed that ASC can not only achieve arbitrary objectives specified by I-Apps, it actually outperformed all prior approaches even in the specific objectives they were designed to achieve. We envision that the framework introduced in this paper will serve as the foundational stepping stone for the research community to build further on it and harness the potential that the learning-based approaches hold for TCP congestion control. While our current implementation is focused on TCP, the underlying mechanisms of ASC can be extended to support other transport protocols such as QUIC, particularly due to QUIC’s user-space stack flexibility. We plan to investigate this in our future work.